OpenStudy (anonymous):

2/3x = 10 a.30 b.15 c.6 1/2 d.6 2/3

OpenStudy (bibby):

you have to isolate x. what could you do to x to isolate it?

OpenStudy (anonymous):

idk , im tired asf its like 1am im tryna get all my work done .

OpenStudy (anonymous):

i dont kno sht

OpenStudy (anonymous):

This is not hard Morgan. You can do this. Start from the bibby's question again. How do you bring x by itself?

OpenStudy (anonymous):

i really have no idea i never pay attention to anything

OpenStudy (hockeychick23):

ok so the easiest way to do this is to multiply both sides by 3/2

OpenStudy (anonymous):

is that 2x/3 = 30?

OpenStudy (bibby):

you have to divide by the thing attached to the x for example, 2x=20 x=20/2=10

OpenStudy (anonymous):

thts so fkn simple

OpenStudy (anonymous):

yes it is but you're tired af.

OpenStudy (bibby):

I am too, I can still do math

OpenStudy (anonymous):

i have A LOT of vocabulary to do , can anyone help me with tht ??

OpenStudy (bibby):

google has a built-in dictionary

OpenStudy (anonymous):

LOOK AT THE TIME. PHEW

OpenStudy (bibby):

TIME TO GET A WATCH. HAH.

OpenStudy (anonymous):

wat about gum ?

OpenStudy (anonymous):

can yall help me in GUM ?